### v1

@@ 5 Connecting the mathematical and substantive models
For the Russian roulette example, the problem with the deterministic potential outcome model is that it assigns the ultimate outcome under $R_{p}$ as a property of the individual rather than as a product of inherent randomness in the treatment. To make this clearer, consider the Russian roulette treatment as including the spinning of the cylinder as well as the firing of the shot: the assumption is that the outcome of the spin is random and is not predictable by any characteristics of the individual. To justify the probabilistic model, one could imagine the spinning to be performed by a machine under the control of an external random number generator.
@@
Another way to see this is to consider an analogous trolley problem, in which a runaway trolley is heading toward a crowd and, if nothing is done, it is expected to kill 1/6 of the people there, but there is an option to switch it to an alternative track leading to a less densely-populated zone where it would be expected to kill only 1/7 of the people in the crowd. Under classical utility theory, switching is an unambiguously dominant strategy. But under the asymmetric utility function (1), if we define deterministic potential outcomes—that is, if we condition on the set of people who would die under the different alternatives—it would be better not to switch.
@@
Or, to morph this into a medical problem, suppose that for a certain otherwise-fatal snakebite the current treatment is an antidote which works perfectly except for 1/6 of the population who have a certain genetic condition. An alternative antidote is proposed which works perfectly except for 1/7 of the population who have a different genetic condition whose occurrence is statistically independent of the other condition in the population. Further suppose that only one antidote can be given to all patients—or, equivalently, that we are treating one patient whose genetic conditions are unknown, and only one antidote can be tried. The decision of whether to switch to the second treatment is equivalent in expected value to the earlier-considered decision to switch from $R_{1/6}$ to $R_{1/7}$, but it corresponds to a different distribution of principal strata and thus a different decision recommendation when using the asymmetric utility function that depends on unrealized outcomes.
When mapping this utility function to a decision recommendation, it matters where the conditioning is done. We see an analogy to the Heisenberg principle in physics, under which outcomes can depend on the order of measurement. In classical probability (in physics jargon, “Boltzmann statistics”), unmeasured data can be expressed as part of a joint distribution, and the order in which information is revealed does not affect observables. In quantum probability, indeterminacy manifests as superposition of wave functions, not averaging over probability, which is why we see the characteristic interference pattern of a two-slit experiment, for example.
@@
The argument, then, is that if you want to use this sort of an asymmetric utility function, you cannot simply integrate out the uncertainty of stochastic potential outcomes ahead of time. The difference between the Russian roulette problem and the snakebite problem is that, in the latter, the potential outcomes are latent variables that are deterministic attributes of the individuals.
@@
How would this apply in real-world outcomes? We consider two areas: medicine and business.
@@
  • In a cleanly-specified medical example in which the outcome is binary (survival or death), when would it make sense to prefer to stick with a treatment that leaves more people dead? There is a vast literature on the trolley problem, and the resolution often turns on issues of agency. One rationale for preferring the status quo, even if the new treatment increases the expected number of lives saved, is that this sort of “social engineering” would reduce people’s control over their own lives. In the Russian roulette example, this reasoning would not apply because the decision to switch from $R_{1/6}$ to $R_{1/7}$ does not harm any identifiable patient.
@@
  • When making a business decision—for which the negative outcome for an individual is not death but rather is to no longer be a customer—there are various reasons to prefer the status quo even if it is stochastically dominated by an alternative treatment. Making changes (“churn”) has its own costs, and one can easily imagine a scenario in which a loss of 1/7 of a customer base, even if accompanied by a gain of 1/6, would represent a net loss. This would be the trolley or snakebite scenario. In the Russian roulette scenario, the decision to switch from $R_{1/6}$ to $R_{1/7}$ would simply result in a reducing the loss by 1/42, which is unambiguously better, as there are no particular customers who could be identified as being lost under one treatment or the other.
@@
To push this further, consider a setting where the potential outcomes are mathematically deterministic but effectively random. For example, suppose that for some reason a company is about to lose all customers whose Social Security numbers are divisible by 6, and an intervention is proposed by which the company will instead lose all customers whose Social Security numbers are divisible by 7. Here the potential outcome is deterministic and we are in the trolley or snakebite scenario, but for all practical purposes this seems more like Russian roulette, in that there are no relevant distinguishing characteristics of people in the four principal strata defined by the deterministic outcomes $(y^{0},y^{1})$. In this case it would seem to make sense to integrate over this uncertainty, that is, to model the potential outcomes as stochastic. What this last twist on our example demonstrates is that the appropriate parameterization ultimately depends not just on the asymmetric utility function but also on its underlying motivation. In real-world settings, we would expect the units in the principal strata to differ in relevant ways, and we expect it would make sense to employ a stochastic potential-outcome model where the distribution for $(y^{0},y^{1})$ depends on observed and latent characteristics of the experimental units.
@@
For another example, imagine there are two kinds of patients presenting with headaches. For one type of patient, the headache is caused by a migraine and is not responsive to treatment; for the other the headaches do respond to treatment with some probability of success. The potential outcome is deterministic for the migraine patients but, depending on how much is known about the mechanism of the treatment, could be stochastic for the non-migraine patients. This parameterization of potential outcomes would make sense even if individual patients’ migraine status is unknown, because the possible action of the treatment depends on characteristics of the individuals. When comparing to a new treatment that works for some migraine patients but at the cost of reduced efficacy for others, one might want to use an asymmetric utility function that depends on unrealized potential outcomes, which in this case would map to migraine status.
@@ 6 Discussion
There is a large literature on decision rules and utility functions that go beyond classical decision theory. There could be a general conservatism or disutility of making changes, just because churn has its own costs—or, in other settings, novelty could be considered to having an inherent benefit as a counter to status quo bias. Preferences can also be based on agency: people can be more comfortable with risks that they feel they have the power to control and can resist an apparently clear improvement if it represents a perceived loss of power or status. And meta-decision issues come up, ideas such as satisficing (Simon, 1958), Type II rationality (Good, 1971), and fast and frugal heuristics (Gigerenzer and Goldstein, 1996) that recognize the cost of gathering and evaluating the information required to make theoretically optimal decisions. There is an expression in poker, “Evaluate the strategy, not the play,” and one might choose relatively inefficient decision heuristics as part of a global decision-making strategy that is cheaper (less costly to evaluate, exerts a lower cognitive or moral load, etc.). One could imagine a “First, do no harm” principle being morally and socially valuable, to the extent that it could be worth making some objectively bad decisions so as to preserve a larger principle of consistently, in the same way that we like to say we live in “a government of laws, not of men,” and so there will be unjust decisions that still need to be enforced in order to not lose that consistency.
The relevance of these ideas to the present paper is that, once we go beyond classical (Neumann-Morgenstern) decision theory and allow the utility of a decision to depend on something other than its outcomes (for example, life or death in the simple medical or Russian roulette example), stochastic potential outcomes can no longer be thought of as equivalent to population averages of discrete potential outcomes. This is a good thing, in that it allows us to apply the asymmetric utility function of Ben-Michael et al. (2024) and Christy and Kowalski (2024) to the trolley or snakebite scenario without having to swallow the uncomfortable conclusion that it is better not to switch from $R_{1/6}$ to $R_{/17}$ in the Russian roulette example. That is, we can attack those those different problems with the same utility theory, just expressing their potential outcomes in a different way.

### v2

@@ 5 Connecting the mathematical and substantive models
We consider examples to explore the role of uncertainty within and variation between units. For the Russian roulette example, the problem with the deterministic potential outcome model is that it assigns the ultimate outcome under $R_{p}$ as a property of the individual rather than as a product of inherent randomness in the treatment. To justify the probabilistic model, one could imagine the spinning to be performed under the control of an external random number generator.
@@
To frame it as a medical problem, suppose that for a certain otherwise-fatal snakebite the current treatment is an antidote that works perfectly except for 1/6 of the population who have a certain genetic condition. A proposed alternative works perfectly except for 1/7 of the population who have a different condition whose occurrence is statistically independent of the other condition in the population. Further, suppose that only one antidote can be given to all patients—or, equivalently, that we are treating one patient whose genetic conditions are unknown, and only one antidote can be tried. The decision of whether to switch to the second treatment is equivalent in expected value to the earlier-considered decision to switch from $R_{1/6}$ to $R_{1/7}$, but it corresponds to variation across units rather than uncertainty within units and thus a different decision recommendation when using the asymmetric utility function that depends on unrealized outcomes.
@@
When mapping this utility function to a decision recommendation, it matters where the conditioning is done. If you want to use this sort of asymmetric utility function, you cannot simply integrate out the uncertainty of stochastic potential outcomes ahead of time. The difference between the Russian roulette problem and the snakebite problem is that, in the latter, the potential outcomes are latent variables that are deterministic attributes of the individuals.
@@
How would this apply to real-world outcomes? We consider two areas: medicine and business.
  1. 1.
@@
    In a cleanly specified medical example in which the outcome is binary (survival or death), when would it make sense to prefer to stick with a treatment that leaves more people dead? There is a vast literature on the trolley problem, and the resolution often turns on issues of agency. One rationale for preferring the status quo, even if the new treatment increases the expected number of lives saved, is that this sort of “social engineering” would reduce people’s control over their own lives. In the Russian roulette example, this reasoning would not apply because the decision to switch from $R_{1/6}$ to $R_{1/7}$ does not harm any identifiable patient.
  2. 2.
@@
    When making a business decision for which a person’s negative outcome is not death but rather is to no longer be a customer, there are various reasons to prefer the status quo even if it is stochastically dominated by an alternative treatment. Making changes (“churn”) has its own costs, and one can easily imagine a scenario in which a loss of 1/7 of a customer base, even if accompanied by a gain of 1/6, would represent a net loss. This would be the trolley or snakebite scenario. In the Russian roulette scenario, the decision to switch from $R_{1/6}$ to $R_{1/7}$ would simply result in a reducing the loss by 1/42, which is unambiguously better, as there are no particular customers who could be identified as being lost under one treatment or the other.
@@
To push this further, consider a setting where the potential outcomes are mathematically deterministic but effectively random. For example, suppose that for some reason a company is about to lose all customers whose Social Security numbers are divisible by 6, and an intervention is proposed by which the company will instead lose all customers whose Social Security numbers are divisible by 7. Here the potential outcome is deterministic as in the snakebite scenario, but it is also similar to Russian roulette, in that there are no relevant distinguishing characteristics of people in the four principal strata defined by the deterministic outcomes $(y^{0},y^{1})$. In this case it could make sense to integrate over this uncertainty, that is, to model the potential outcomes as stochastic. What this last twist on our example demonstrates is that the appropriate parameterization ultimately depends not just on the asymmetric utility function but also on its underlying motivation. In real-world settings, we would expect the units in the principal strata to differ in relevant ways, and we expect it would make sense to employ a stochastic potential-outcome model where the distribution for $(y^{0},y^{1})$ depends on observed and latent characteristics of the experimental units.
@@
We have so far considered the two extreme cases in which variation arises either all from within-unit uncertainty (as with the Russian roulette example) or all from between-unit variation (as with the snakebite example). We will now consider an intermediate example in which both levels of variation are relevant. Imagine there are two kinds of patients presenting with headaches. For one type of patient, the headache is a migraine that is not responsive to treatment; for the other the headaches do respond to treatment with some probability of success. The potential outcome is deterministic for the migraine patients but, depending on how much is known about the mechanism of the treatment, could be stochastic for the non-migraine patients. This parameterization of potential outcomes would make sense even if individual patients’ migraine status is unknown, because the possible action of the treatment depends on characteristics of the individuals. When comparing to a new treatment that works for some migraine patients but at the cost of reduced efficacy for others, one might want to use an asymmetric utility function that depends on unrealized potential outcomes. Using stochastic potential outcomes to account for the within-unit uncertainty for the non-migraine patients allows us to avoid the inconsistent policy recommendations we encountered in the Russian roulette example, while maintaining the benefits of a “first do no harm” policy on the level of across-unit variability.
@@ 6 Discussion
There is a large literature on decision rules and utility functions that go beyond classical decision theory. There could be a general conservatism or disutility of making changes, just because churn has its own costs—or, in other settings, novelty could be considered to having an inherent benefit as a counter to status quo bias. Preferences can also be based on agency: people can be more comfortable with risks that they feel they have the power to control and can resist an apparently clear improvement if it represents a perceived loss of power or status. And meta-decision issues come up, ideas such as satisficing [22], Type II rationality [10], and fast and frugal heuristics [9] that recognize the cost of gathering and evaluating the information required to make theoretically optimal decisions. There is an expression in poker, “Evaluate the strategy, not the play,” and one might choose relatively inefficient decision heuristics as part of a global decision-making strategy that is cheaper (less costly to evaluate, exerts a lower cognitive or moral load, etc.). One could imagine a “First, do no harm” principle being morally and socially valuable, to the extent that it could be worth making some objectively bad decisions so as to preserve a larger principle of consistently, in the same way that we like to say we live in “a government of laws, not of men,” and so there will be unjust decisions that still need to be enforced in order to not lose that consistency.
The relevance of these ideas to the present paper is that, once we go beyond classical (Neumann-Morgenstern) decision theory and allow the utility of a decision to depend on something other than its outcomes (for example, life or death in the simple medical or Russian roulette example), stochastic potential outcomes can no longer be thought of as equivalent to population averages of discrete potential outcomes. This is a good thing, in that it allows us to apply the asymmetric utility function of [2] and [3] to the trolley or snakebite scenario without having to swallow the uncomfortable conclusion that it is better not to switch from $R_{1/6}$ to $R_{/17}$ in the Russian roulette example. That is, we can attack those those different problems with the same utility theory, just expressing their potential outcomes differently.
@@
Acknowledgement
We thank Amanda Kowalski, the editor, and two reviewers for helpful discussions and the U.S. Office of Naval Research for partial support.